Lakeview is 69 miles from Springfield. Irania must go to the library in Springfield to get a document for the law firm for which she works. Find how long it takes for her to drive round-trip if she averages 43 mph.
Answer by Alan3354(69443)

The round trip is 69*2 = 138 miles
---------
d = r*t
t = d/r